# library.bib

### v1

@@ Methods
Table 1: Performance metrics for predicting $T$ from $pki$ with UNet and DDUNet. Errors in [°C], MSE in [°C2], PAT in [%], and SSIM unitless. Inferred on the test of 3dp and the scaling datapoint.
Model Test Scaling
L MAE MSE PAT SSIM L MAE MSE PAT SSIM
UNet3dp 4.8642 0.1314 0.0492 39.05 0.5794
UNet101dp 4.3985 0.0473 0.0100 13.63 0.9827 4.3426 0.0202 0.0033 4.33 0.9955
$2\times 2$-DDUNet101dp 3.4257 0.0548 0.0128 17.42 0.9804 4.1806 0.0235 0.0052 6.11 0.9940
@@ 5 Methodology
In data-scarce settings, performance can improve when the structure of a model reflects the underlying physics, cf. Section 2. This section hence outlines the relevant physical processes, motivates our LGCNN method, and discusses a few details of LGCNN’s individual components.
@@
Transport of heat induced by heat injections of GWHPs in the subsurface is an advection–diffusion process. Since the Péclet number, i.e., the advection-diffusion ratio, is high in our application (Pemuch-$\gg$1, more in Section A.1), the system is advection-dominated. Advection is largely governed by the global hydraulic pressure gradient, driving flow from higher to lower regions. Locally, the flow paths are influenced by the spatial distribution of permeability $k$, cf. Figure 2. As a result, small changes in $k$ can lead to significant differences in flow paths further downstream, demonstrating high input sensitivity.
@@
In a fully coupled physical model, flow field computation and heat transport along streamlines [streamline] starting at heat pumps (advection) and their widening (diffusion) must be solved monolithically. By splitting effects into local and global and working with steady-state simulated $\vec{v}$ during training, we can reduce the coupling to one direction without significant loss in accuracy. Our simplified physical pipeline consists of three steps: (1) compute a steady-state flow field $\vec{v}$ from initial subsurface parameters $p$ and $k$, with $i$ encoding mass influx around GWHPs; (2) transport injected heat along streamlines governed by $\vec{v}$ until quasi-steady state[bawue_arbeitshilfe]; and (3) apply plume widening to these heat paths, informed by soil diffusivity and $\vec{v}$, to approximate diffusion effects. The resulting one-way coupled LGCNN approach (see Figure 1) can formally be described as:
  1. Step 1
    Velocities (local) We employ a CNN to predict the velocity field $\vec{v}=\left(v_{x},v_{y}\right)$ from $p$, $k$, and $i$:
    $\text{CNN}\left(p,k,i\right)=\vec{v}.$ (1)
  2. Step 2
    Streamlines (global) Based on the predicted velocities $\vec{v}$ from Step 1, we compute streamlines $\vec{s}$ originating from all pump locations in $i$ with an initial value problem (IVP) solver:
    $\text{IVP}\left(i,\vec{v}\right)=\vec{s}.$ (2)
  3. Step 3
    Temperature (local) A second CNN predicts the temperature field $T$ from the initial inputs $p$, $k$, $i$, the predicted velocities $\vec{v}$ of Step 1 and streamlines $\vec{s}$ calculated in Step 2:
    $\text{CNN}\left(p,k,i,\vec{v},\vec{s}\right)=T.$ (3)
The model outputs the steady-state temperature field $T$ and, using the same inputs $p$, $k$, and $i$ as a simulation, serves as its direct surrogate.
@@ 3. Step 3
The local Steps 1 and 3 are approximated by a UNet[CNN15, pelzer2024], see Section A.2 for details of our architecture. It is beneficial to omit zero-padding to enforce shift invariance and reliance on purely local information; this approach is also validated by Islam et al. in [Islam2020How]. Due to the input sensitivity of the entire problem, each step is trained independently on simulated input data. In particular, the CNN in Step 3 is trained on the simulated velocities $\vec{v}_{\text{sim}}$ and streamlines based on these fields. Only during inference, all steps are applied sequentially, i.e., Step 3 uses the outputs of Steps 1 and 2, including the predicted $\vec{v}_{\text{pred}}$ as inputs.
The training is accelerated by partitioning the data into overlapping patches, which increases the effective number of datapoints while reducing their spatial size. Combined with the streamlines’ embedding and localizing of global flow patterns for the fully convolutional (locally) acting CNN, this training data enrichment allows the model to train effectively on very little data, sometimes as little as a single simulation run, while generalizing to unseen and even to larger domains. Validation and testing are conducted on independent simulation runs, with the full domain being processed as a whole instead of patches during inference.
The architecture, input choices, and training hyperparameters — including the patch size and overlap in the datapoint partitioning — were optimized through a combination of manual and automated tree-based search using Optuna [optuna_2019]; see Section A.2.1 for details.
@@
with a lightweight numerical solver for each heat pump, where $y_{0}$ represents the location of a heat pump in $i$. We employ the solve_ivp function from SciPy [scipy] with an implicit fifth-order Runge-Kutta method [scipyRadaux2014] and 10 0001000010\,00010 000 time steps of one day, corresponding to the simulation time of forming a heat plume.
The computed streamlines $y$ represent a sequence of positions in 2D, that a particle - injected at a heat pump’s location - would traverse based on $\vec{v}$. The corresponding grid cells are assigned values that fade linearly from 1 to 0, reflecting the time it takes to reach each position, similar to a soft one-hot encoding.
To address the sensitivity of this process with respect to $k$ and $\vec{v}$, we calculate a set of outer streamlines $s_{o}$. For this, we perturb each heat pump’s location $y_{0}$ by 10 cells orthogonal to the global flow direction, i.e., $\nabla p$, and compute the corresponding streamlines. By incorporating $\vec{s}=\left(s,s_{o}\right)$, the subsequent Step 3 obtains a spatial, localized representation of flow paths. In future work, we aim to explore probabilistic perturbations to compute the mean and standard deviation of streamlines, which, while being computationally more expensive, could be efficiently parallelized on GPUs.
6 Results on Synthetic Permeability Fields
This section evaluates our model’s performance on the baseline dataset 3dp and on a scaling test datapoint; cf. Section 3. In Section 6.1, we present metrics for isolated velocity (Step 1) and temperature (Step 3) predictions, and for the full pipeline, demonstrating the potential of our approach. In ablation studies (Section 6.2), we motivate model design choices. Additional metrics and visualizations can be found in Section A.3.
Table 2: Metrics for predicting $\vec{v}_{\text{pred}}$ in Step 1 and in ablation studies of Step 1 on test and on scaling data from our baseline 3dp dataset with random $k$. In [m/y], except for MSE in [m2/y2], SSIM unitless.
Output Case L MAE MSE SSIM
LGCNN Step 1
$v_{x}$ test 190.8046 22.3178 972.5668 0.9911
$v_{y}$ test 256.2519 32.7444 2031.3357 0.9812
$v_{x}$ scaling 294.0457 24.9261 1204.1154 0.9911
$v_{y}$ scaling 367.6891 26.2795 1463.8218 0.9820
Ablation study Train without data partitioning
$v_{x}$ test 202.6397 23.8584 1144.4479 0.9913
$v_{y}$ test 332.9557 35.9249 2490.4104 0.9789
[…]
$v_{x}$ test 1047.8464 12.7036 896.2758 0.9016
$v_{y}$ test 2594.9504 14.3878 2317.6863 0.9903
$v_{x}$ scaling 659.8439 5.2152 79.2508 0.9699
$v_{y}$ scaling 456.1904 5.7492 83.9735 0.9948
[…]
$v_{x}$ test 2453.5881 13.6735 1367.9061 0.9325
$v_{y}$ test 2029.0269 15.1332 2221.4226 0.9919
$v_{x}$ scaling 421.4186 8.1184 125.9511 0.9314
$v_{y}$ scaling 400.2490 7.4200 135.7929 0.9955
Table 3: Metrics for predicting $T$ in Step 3, in the full pipeline, and in ablations studies of Step 3 on test and on scaling data from our baseline 3dp dataset with random $k$. In [°C], except for MSE in [°C2], PAT in [%], SSIM unitless.
Case L MAE MSE PAT SSIM
LGCNN Step 3, i.e., test on $\vec{v}_{\text{sim}}$
test 2.8990 0.0347 0.0041 7.54 0.9304
scaling 3.0250 0.0168 0.0014 2.05 0.9510
Pipeline, i.e., test on $\vec{v}_{\text{pred}}$
test 4.2120 0.0905 0.0307 28.92 0.7637
scaling 4.9366 0.0413 0.0141 10.87 0.8654
Ablation study Train in sequence
test 4.0804 0.0901 0.0289 29.44 0.7553
Include zero-padding
test 2.3391 0.0418 0.0048 9.19 0.8838
Train without data partitioning
test 3.6420 0.0351 0.0043 7.41 0.9245
[…]
test 4.4658 0.0416 0.0067 11.69 0.9835
scaling 4.3489 0.0176 0.0022 3.07 0.9954
[…]
test 4.2772 0.0376 0.0058 10.96 0.9915
scaling 3.9946 0.0192 0.0025 4.73 0.9968
6.1 Performance of LGCNN
@@
Performance metrics for Step 1 are shown in Table 3, for Step 3 (trained and tested on simulated $\vec{v}_{\text{sim}}$) in Table 3. The combination of quantitative (MAE of (22.3, 32.7) m/y and 0.03 °C) and qualitative results through a visual assessment of Figure 3 demonstrate a strong performance with heat plumes forming according to $k$.
Figure 3: $pki\vec{s}\vec{v}_{\text{sim}}\rightarrow T$ (baseline dataset (3dp), test). 1st column: Label $T$. 2nd-3rd column: LGCNN Step 3, trained without and with zero-padding.
@@
Testing the full pipeline, i.e., Steps 2 and 3 are based on $\vec{v}_{\text{pred}}$, results in a higher test MAE of 0.09 °C compared to an isolated Step 3 (see Table 3). This was expected as the isolated Step 3 uses the true, simulated velocity fields $\vec{v}_{\text{sim}}$ as inputs. Visual assessment of representative predictions in Figure 2 reveals physically plausible heat plumes in terms of shape, extent, and heat magnitude. Deviations in streamlines arise from smaller errors in the velocity predictions from Step 1 $\vec{v}_{\text{pred}}$, highlighting the input sensitivity, especially near bifurcations, where small perturbations in $\vec{v}$ can lead to an alternative path around a clay lens in $k$. Compared to fully data-driven models trained on 101dp (see Table 1), our model achieves a similar accuracy while training on only 1 versus 73 training datapoints, which strongly reduces computation time (see Table 16 for details).
Scaling Test
We can already expect that our model is able to scale to larger domains, as we train only on local patches of the spatial domain and achieve high test accuracies on the full domain during validation and testing; see also Section 6.2 Training Data Partitioning for more details. To further demonstrate the scaling, we test our model on a domain of 4$\times$ the size of the training domain with the same number of heat pumps, hence with a lower density of GWHPs/km2; cf. Figure 4.
[…]
We obtain a comparable MAE of 0.02 °C for Step 3, and an MAE of 0.04 °C for the full pipeline, see Table 3. A visual assessment of Figure 4 yields similar qualitative behavior as the previous results in Figure 2.
6.2 Ablation Study
Training in Sequence
[…]
Zero-Padding
Removing zero-padding from all convolutional layers yields smaller but cleaner output fields, as shown in Figure 3. As a result, we obtain improved performance, reflected in lower PAT and higher SSIM scores in Table 3.
Partitioning Training Data
Hyperparameter search shows the best performance for partitioning the domain during training into 20 736 overlapping patches of 256$\times$256 cells for Step 1 and 82 944 patches for Step 3. Although this increases per-epoch compute time compared to full-domain training, it significantly reduces the number of epochs needed to reach comparable test performance (cf. Tables 3 and 3), resulting in an overall training time reduction of 67–90% (cf. Table 15).
[…]
We individually replace each of the three steps with an optimized (cf. Section A.2.1) vanilla approach, i.e., UNet or DDUNet trained on 101dp. For Step 2, both fail completely; results for Steps 1 and 3 are summarized in Tables 3 and 3. In Step 1, UNet slightly outperforms DDUNet, while the reverse holds for Step 3. A visual comparison of the scaling datapoint in Figure 4 reveals that increasing training data from 3dp to 101dp has minimal impact on the performance in Step 3.
Modifying Streamline Inputs of Step 3
We evaluate the influence of Step 3’s inputs, focusing on the effect of modifying or omitting individual components of the streamlines $\vec{s}=\left(s,s_{o}\right)$. When both the central streamlines $s$ and offset streamlines $s_{o}$ are excluded, predicted plumes fail to follow flow paths beyond the CNN’s receptive field. Omitting only $s_{o}$ results in overly narrow plumes, whereas excluding only $s$ produces overly smeared temperature fields. Including both but removing the fading of values along the streamlines leads to diffuse background temperatures and less localized plumes. Overall, the Huber validation loss increases by 32–132%, depending on the experiment. For visual, quantitative, and extended results, we refer to Section A.3.
7 Domain Transfer to Real Permeability Fields
Finally, we investigate whether our approach transfers to the realistic dataset in Section 3, which is based on real-world permeability fields $k$ that are cut from geological maps of Munich. Since these maps cover a limited area, the number of potential datapoints is very small. We compare performance for temperature prediction of Step 3 and the full pipeline on the validation and scaling datapoints. Results are presented in Tables 5 and 5; additional metrics are in Section A.3.
Table 4: Metrics for predicting $\vec{v}_{\text{pred}}$ in Step 1 on validation and scaling data from our realistic dataset. In [m/y], except for MSE in [m2/y2], SSIM unitless.
Output Case L MAE MSE SSIM
Step 1
$v_{x}$ val 106.8607 15.4095 380.3762 0.9939
$v_{y}$ val 74.5570 10.6605 148.2475 0.9993
$v_{x}$ scaling 459.1620 110.0078 13570.0000 0.9462
$v_{y}$ scaling 240.9406 17.6051 616.5186 0.9965
Table 5: Metrics for predicting $T$ in Step 3 and in the full pipeline on validation and scaling data from our realistic dataset. In [°C], except for MSE in [°C2], PAT in [%], SSIM unitless.
Case L MAE MSE PAT SSIM
Step 3, i.e., test on $\vec{v}_{\text{sim}}$
val 0.8222 0.0175 1.01e-3 2.18 0.9672
scaling 0.8052 0.0189 8.21e-4 0.92 0.9497
Full pipeline, i.e., test on $\vec{v}_{\text{pred}}$
val 2.3194 0.0841 2.75e-2 27.79 0.7510
scaling 2.0511 0.0394 4.42e-3 10.02 0.8708
Methodological Adaptations
Slight adjustments to chunk size, overlap, data split, and architecture yield better performance compared to those used on the baseline dataset, see Section A.2.1 for details. Visual inspection of input $k$ in Figure 5 reveals fewer but larger-scale features. Therefore, training can benefit from a larger spatial context, which requires more training data. We increase the number of training samples by using three out of the four available datapoints for training and one for validation. For the streamlines computation, we switch to an explicit fourth-order Runge–Kutta scheme, which proved stable and sufficient for training and inference due to the lower frequency in $k$.
Performance and Scaling Tests
In Step 3, we reach an even lower maximum absolute error $L_{\infty}$ of 0.82 °C compared to 2.90 °C of our initial model on the baseline dataset (Section 6). Other metrics confirm a similar performance; see Table 3. In the scaling test, most errors are comparable too. For the full pipeline, errors are slightly higher on the validation and scaling data compared to an isolated Step 3, consistent with our observations on the baseline dataset (Table 3). Physical consistency, which is our primary objective, remains strong: Shape, magnitude, and connectivity of the predicted heat plumes are well preserved, even in the scaling test (Figure 5). Local deviations are primarily due to differences between $\vec{v}_{\text{sim}}$ and $\vec{v}_{\text{pred}}$, as evident from the discrepancy between Step 3 and full pipeline outputs.
Figure 5: $pki\rightarrow T$ (realistic dataset, scaling). 1st & 2nd column: Label $T$ and input $k$. 3rd-4th column: LGCNN (isolated Step 3 and full pipeline).
8 Conclusions and Future Work
In real-world scientific applications, data is often very limited. At the same time, models are expected to generalize and scale to large domain sizes, pushing standard data-driven models to their limits. To address this issue, we proposed the LGCNN model, a hybrid, physics-inspired CNN approach. In the prediction of groundwater temperature fields under arbitrary heat pump configurations and real permeability data, we identify long-range interactions driven by advection as the problematic process for classical deep learning models, such as vanilla CNNs. We replace this part of the prediction with a simple numerical solver, resulting in a sequence (CNN 1 – numerical solver – CNN 2). While this setup is specific to groundwater heat transport, the general approach should be applied to a wider range of applications. Finally, we stress that traditional error metrics can obscure effects like input sensitivity; therefore, we additionally rely on visual assessments and problem-specific metrics such as PAT. In future work, we will address dataset limitations, investigate input sensitivity, and accelerate streamline computation.
Acknowledgements
We thank the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for supporting this work by funding - EXC2075 – 390740016 under Germany’s Excellence Strategy. We acknowledge the support by the Stuttgart Center for Simulation Science (SimTech). We also acknowledge the use of computational resources of the DelftBlue supercomputer, provided by Delft High Performance Computing Centre [DHPC2024]. We thank Kai Zosseder and Fabian Böttcher at the Technical University of Munich for their application-related input.
[…]
Data and Code Accessibility
Our raw data and pretrained models are publicly available on DaRUS https://darus.uni-stuttgart.de/dataverse/ipvs_lgcnn/ or more specifically:
Raw datasets:
Pretrained models (including hyperparameters):
@@ Acknowledgements
  • LGCNN on real permeability fields, 4dp: https://doi.org/10.18419/darus-5082

### v2

@@ 4 Methodology
When generic ML methods fail in data-scarce settings, performance can be improved by adopting a more task-specific approach whose structure reflects the underlying physics (Fernández-Godino et al., 2024; Thodi et al., 2024; Zhang et al., 2020; Yousefzadeh et al., 2025; Bertels & Willems, 2023). Accordingly, this section outlines the relevant physical processes that motivate the proposed LGCNN method and discusses key aspects of its components.
@@
[…]
@@
In a fully coupled simulation, flow field computation and heat transport along streamlines (str, 2011) starting at heat pumps (advection) and their widening (diffusion) influence each other. Therefore, convergence requires that two-way coupling conditions are satisfied, whether using a monolithic or a segregated approach.
By working with steady-state simulated $\vec{v}$ as interim outputs during training (cf. Section˜A.1 Reasoning for Decoupling Assumptions), we can split the processes into local (diffusion) and global (advection) dominated ones and hence extract three subsequent steps that are coupled in only one direction. Our simplified physical pipeline consists of three steps: (1) compute a steady-state flow field $\vec{v}$ from initial subsurface parameters $p$ and $k$, with $i$ encoding mass influx around GWHPs; (2) transport injected heat along streamlines governed by $\vec{v}$ until quasi-steady state (Umweltministerium Baden-Württemberg, 2009); and (3) apply plume widening to these heat paths, informed by soil diffusivity and $\vec{v}$, to approximate diffusion effects. The resulting one-way coupled application of LGCNN (see Figure˜1) can formally be described as:
  1. Step 1:
    Velocities (local) We employ a CNN to predict the velocity field $\vec{v}=(v_{x},v_{y})$ from $p$, $k$, and $i$:
    $\text{CNN}(p,k,i)=\vec{v}.$ (1)
  2. Step 2:
    Streamlines (global) Based on the predicted velocities $\vec{v}$, we compute streamlines $\vec{s}$ originating from all pumps in $i$ with an initial value problem (IVP) solver:
    $\text{IVP}(i,\vec{v})=\vec{s}.$ (2)
  3. Step 3:
    Temperature (local) A second CNN predicts the temperature field $T$:
    $\text{CNN}(k,i,\vec{v},\vec{s})=T.$ (3)
The model outputs a steady-state temperature field $T$ using the same inputs as a simulation, hence serving as its direct surrogate.
@@ 3. Step 3:
The local Steps 1 and 3 are approximated by a UNet, same as in other groundwater applications (Davis et al., 2023; Pelzer & Schulte, 2024), see Section˜A.2.1 for details of our architecture. It is beneficial to omit zero-padding to enforce shift invariance and reliance on purely local information (Islam et al., 2020). Due to the input sensitivity of the entire problem, the CNN in Step 3 is trained on simulated velocities $\vec{v}_{\text{sim}}$ and streamlines based on these fields. Only during inference, all steps are applied sequentially, i.e., Step 3 uses the outputs of Steps 1 and 2, e.g., by including the predicted $\vec{v}_{\text{pred}}$ as inputs.
The training is accelerated by partitioning the data into overlapping patches, which increases the effective number of datapoints while reducing their spatial size. Combined with the streamlines’ embedding and localizing of global flow patterns for the fully convolutional (locally) acting CNN, this training data enrichment allows the model to train effectively on very little data, partially as little as a single simulation run, while generalizing to unseen and even to larger domains. Validation and testing are conducted on independent simulation runs, with the full domain being processed as a whole instead of patches during inference.
The architecture, input choices, and training hyperparameters, including patch size and overlap in the datapoint partitioning, were optimized; see Section˜A.2.3 for details.
@@
with a lightweight numerical solver for each heat pump, where $y_{0}$ represents the location of a heat pump in $i$. We employ an implicit fifth-order Runge–Kutta method (sci, ) with 10,000 one-day time steps, corresponding to the physical timescale of heat plume formation.
The computed streamlines $y$ represent sequences of 2D positions traversed by a particle injected at $y_{0}$ under the velocity field $\vec{v}$. These paths are embedded into 2D fields along the traversed grid cells by assigning values that decay linearly from 1 to 0, reflecting the time required to reach each cell and yielding a time-weighted 2D-embedding.
To address the sensitivity of this process with respect to $k$ and $\vec{v}$, we calculate the outer streamlines $s_{o}$. For this, we perturb each heat pump’s location $y_{0}$ by $\pm$10 cells orthogonal to the global flow direction, i.e., $\nabla p$, and compute the corresponding streamlines. By incorporating $\vec{s}=\left(s,s_{o}\right)$ as two inputs to Step 3, it obtains a spatial, localized representation of flow paths. In future work, we aim to explore probabilistic perturbations to compute the mean and standard deviation of streamlines, which, while being computationally more expensive, could be efficiently parallelized on GPUs.
5 Results
We evaluate LGCNN’s performance on the baseline dataset 3dp for isolated Steps 1 and 3, and for the full pipeline (end-to-end), demonstrating the potential and scalability of our approach. In Section˜5.2, we motivate model design choices through variational experiments. Additional metrics and visualizations can be found in Section˜A.3. Section˜5.3 explains adaptations, performance, and scalability on the more realistic dataset.
5.1 Performance on the Synthetic Dataset
@@
Performance metrics in Table˜1 Ours are for Step 3 (trained and tested on simulated groundtruth $\vec{v}_{\text{sim}}$), and for Step 1 highlighted in gray. The combination of quantitative (MAE of (22.3, 32.7) m/y and 0.04 °C, and only 0.07% unconnected cells) and qualitative measures (visual assessment of Figure˜3) demonstrate a strong, physically reasonable performance, especially compared to a generic UNet trained on the same number of datapoints (Figure˜2).
Figure 3: Step 3: $pki\vec{s}\vec{v}_{\text{sim}}\rightarrow T$ (baseline dataset (3dp): test). 1st column: Label $T$. 2nd-3rd column: LGCNN Step 3, trained without and with zero-padding.
@@
Testing the full pipeline, i.e., Steps 2 and 3 applied to $\vec{v}_{\text{pred}}$ predicted by Step 1, results in a higher test MAE of 0.09 °C compared to an isolated Step 3. This was expected as the isolated Step 3 uses groundtruth, simulated velocity fields $\vec{v}_{\text{sim}}$ as inputs. Visual assessment of representative predictions in Figure˜2 reveals physically plausible heat plumes in terms of shape, extent, and heat magnitude, also measurable by a comparably low MoC of 0.09 %. Deviations in streamlines arise from smaller errors in the velocity predictions $\vec{v}_{\text{pred}}$ from Step 1, highlighting the input sensitivity of the physical problem, especially near bifurcations, where small perturbations in $\vec{v}$ can lead to an alternative path, e.g., around a clay lens in $k$. Compared to training end-to-end on 101dp (see Table˜1 Standard), our model achieves a similar MoC and at least half the accuracy while training on only 1 versus 73 training datapoints, which strongly reduces computation time (see Table˜12 for details) and data requirements.
Scaling
@@
LGCNN obtains a low MAE of 0.02 °C for Step 3, and 0.04 °C for the full pipeline and a very low MoC of 0.06%, see Table˜1. A visual assessment of Figure˜4 yields similar qualitative, physically reasonable behavior as the previous results in Figure˜2, demonstrating the scalability of LGCNN.
5.2 Experiments on the Synthetic Dataset
Training in Sequence
When training Step 3, we consider two alternatives for the velocity input: simulated velocities $\vec{v}_{\text{sim}}$ and velocities predicted by Step 1 $\vec{v}_{\text{pred}}$. Training the model sequentially using $\vec{v}_{\text{pred}}$ results in substantially higher MoC values compared to training with $\vec{v}_{\text{sim}}$, as shown in Table˜1 (Experiments vs. Ours). As illustrated in Figure˜2, using $\vec{v}_{\text{pred}}$ during training leads to physically implausible temperature fields characterized by noisy artifacts and fragmented heat plumes. We attribute this behavior to local misalignments between the predicted streamline inputs $\vec{s}$ and the corresponding temperature label $T$, which strongly complicates localized learning. In contrast, training with $\vec{v}_{\text{sim}}$ produces streamline fields that are physically consistent with the temperature distribution which provides informative gradients and enables stable and effective training.
Enabling Zero-Padding
Using no padding in all convolutional layers produces smaller but cleaner output fields, as shown in Figure˜3. This leads to a slight improvement in standard ML metrics and a substantial improvement in the MoC score (Table˜1, Ours vs. Experiments), indicating fewer disconnected heat plumes.
Partitioning Training Data
Hyperparameter search shows the best performance for partitioning the domain during training into 20 736 overlapping patches of $\times256$ cells for Step 1 and 82 944 patches for Step 3. Although this increases per-epoch compute time compared to full-domain training, it significantly reduces the number of epochs needed to reach comparable test performance (cf. Table˜1 Experiments vs. Ours, Step 1), resulting in an overall training time reduction of 67–90% (cf. Table˜11) while maintaining comparable accuracy across all metrics.
[…]
We replace each step with an optimized UNet or DDUNet trained on 101dp (cf. Section˜A.2.3). Step 2 fails for both models. Results for Steps 1 and 3 are summarized in Table˜1 Experiments. In Step 1, both models struggle to predict $v_{y}$ due to outliers in training data. In Step 3, performance is comparable to 3dp in standard metrics but yields a substantially higher MoC, with scaling results in Figure˜4 showing minimal benefit from increased training data.
Modifying Streamline Inputs of Step 3
We evaluate the influence of Step 3’s inputs, focusing on the effect of modifying or omitting individual components of the streamlines $\vec{s}=\left(s,s_{o}\right)$. When both the central streamlines $s$ and offset streamlines $s_{o}$ are excluded, predicted plumes fail to follow flow paths beyond the CNN’s receptive field. Omitting only $s_{o}$ results in overly narrow plumes, whereas excluding only $s$ produces overly smeared temperature fields. Including both but removing the fading over time along the streamlines leads to diffuse background temperatures and overly precise and too long plumes. Overall, the Huber validation loss increases by 32-132%, depending on the experiment. For extended and visual results, we refer to Section˜A.3.
5.3 Domain Transfer to Real Permeability Fields
We evaluate our method on real permeability fields $k$ (Section˜2), for which only four datapoints are available. Results for Step 3 and the full pipeline on the validation and scaling datapoints are shown in Table˜1 Domain transfer; additional metrics are provided in Section˜A.3.
Methodological Adaptations
Slight adjustments to chunk size, overlap, data split, and architecture yield better performance compared to those used on the baseline dataset, see Section˜A.2.3 for details. Visual inspection of input $k$ in Figure˜5 reveals fewer but larger-scale features. Therefore, training can benefit from a larger spatial context, which requires more training data. We increase the number of training samples by using three out of the four available datapoints for training and one for validation. For the streamlines computation, we switch to an explicit fourth-order Runge–Kutta scheme, which proved stable and sufficient for training and inference due to the lower frequencies in $k$.
Performance and Scaling Tests
In Step 3, we reach an even lower maximum absolute error $L_{\infty}$ of 0.82 °C compared to 2.90 °C of our initial model on the baseline dataset (Section˜5). Other metrics confirm this except for SSIM and MoC; cf. Table˜1 Domain transfer. In the scaling test, most errors behave as before, as expected. For the full pipeline, errors are slightly higher on the validation and scaling data compared to an isolated Step 3, consistent with our observations on the baseline dataset. Physical consistency, which is our primary objective, remains strong: Shape, magnitude, and connectivity of the predicted heat plumes are well preserved, even in the scaling test (Figure˜5). Local deviations are primarily due to differences between $\vec{v}_{\text{sim}}$ and $\vec{v}_{\text{pred}}$, as evident from the discrepancy between Step 3 and full pipeline outputs.
Figure 5: Predictions of $T$ (realistic dataset (4dp): scaling). Input $k$, label $T$ LGCNN isolated Step 3, and full pipeline.
Based on their poor generalization in data-scarce regimes observed on the baseline dataset, we omit generic ML architectures from the experiments on realistic $k$ fields.
6 Conclusions and Future Work
We introduce the Local-Global Convolutional Neural Network (LGCNN) to address complex, data-scarce learning problems governed by both local and non-local physical interactions. Trained on only a few simulations, LGCNN accurately predicts subsurface temperature fields on much larger two-dimensional heterogeneous domains, while providing an inference-time speed-up of roughly 2,000× compared to classical high-fidelity simulations on the training data domain. By combining translation-invariant CNNs for local processes with a lightweight numerical surrogate for global transport, our method achieves data-efficient, scalable, and physically consistent predictions, outperforming conventional ML models such as UNet variants that struggle to capture long-range dependencies in data-scarce regimes. Although demonstrated for groundwater heat transport, the concept of local-global decomposition can be generalized to a broad class of advection-diffusion systems. Future work will extend the approach to three-dimensional and transient scenarios, investigate sensitivity to input uncertainty, and further accelerate streamline computation.
Acknowledgements
We thank the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for supporting this work by funding - EXC2075 – 390740016 under Germany’s Excellence Strategy. We acknowledge the support by the Stuttgart Center for Simulation Science (SimTech). We also acknowledge the use of computational resources of the DelftBlue supercomputer, provided by Delft High Performance Computing Centre (Delft High Performance Computing Centre, DHPC). We thank Kai Zosseder and Fabian Böttcher at the Technical University of Munich for their application-related input.
Reproducibility
To ensure reproducibility, our code to train and evaluate models, the trained and evaluated models themselves, and the datasets designed for training and testing are included in the supplementary material and will be published after the review process. Our raw data and pretrained models are publicly available on SurfDrive, see below in Supplementary Material.